OpenStudy (anonymous):

Easton simplified the following expression: (x^2y^6z^5)(x^4y^5z^3) If he writes his answer in the form of x^a^y^b^z^c, what is the value of b, the exponent on y?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

First thing write it using standard notation: $$(x^2*y^6*z^5)(x^4*y^5*z^3)$$

OpenStudy (skullpatrol):

To multiply two powers having the same base you add the exponents.

OpenStudy (skullpatrol):

$$y^6= y*y*y*y*y*y$$ $$y^5=y*y*y*y*y$$ $$y^6*y^5= (y*y*y*y*y*y)(y*y*y*y*y)$$

OpenStudy (skullpatrol):

What is the value of b, the exponent on y?

OpenStudy (anonymous):

It dose not tell me what the value of b or the exponent y.

OpenStudy (skullpatrol):

Yes, the question is asking you to calculate that.

OpenStudy (skullpatrol):

To multiply two powers having the same base you add the exponents.

OpenStudy (skullpatrol):

$$6+5=?$$

OpenStudy (skullpatrol):

$$?=b$$ which is the exponent on $$y^b$$

OpenStudy (skullpatrol):

If he writes his answer in the form of $$x^ay^bz^c$$,

OpenStudy (anonymous):

b is the exponent of y

OpenStudy (skullpatrol):

$$(x^2y^6z^5)(x^4y^5z^3)=x^ay^bz^c$$

OpenStudy (skullpatrol):